# From Static to Adaptive Defense: Federated Multi-Agent Deep Reinforcement Learning-Driven Moving Target Defense Against DoS Attacks in UAV Swarm Networks

### v2

@@ Abstract
Index Terms:
Low-altitude networks, unmanned aerial vehicles, denial-of-service attacks, moving target defense, federated multi-agent deep reinforcement learning.
@@ I Introduction
Nevertheless, the widespread adoption of UAV swarms also brings new security challenges [7] to low-altitude networks. Due to their reliance on open wireless links, limited energy and processing capabilities, UAV networks are particularly vulnerable to Denial-of-Service (DoS) attacks [8, 9]. For example, UAVs often operate without robust authentication or traffic filtering mechanisms. Attackers can easily launch attacks by overwhelming low-altitude communication channels or computational resources, leading to service disruptions or UAV disconnection from the swarm [10]. In mission-critical scenarios, even a brief loss of connectivity or control can have catastrophic consequences, including the failure of time-sensitive missions or the crash of drones.
@@ II-A DoS Attacks Detection and Mitigation in UAV Networks
DoS attacks in UAV networks have been extensively studied, with various detection and mitigation techniques proposed. For instance, Fu et al. [19] proposed an IDS for UAV networks, integrating convolutional neural network (CNN) and long short-term memory (LSTM) to achieve high detection accuracy of 94.4% for DoS attacks. Similarly, Hassler et al. [20] employed a method that fuses UAV cyber and physical features, which achieves an accuracy of up to 98.5% for attacks. To overcome challenges such as small sample sizes and uneven data distribution among UAVs, He et al. [21] leveraged a conditional generative adversarial network (CGAN)-based intrusion detection algorithm, achieving an accuracy of 99% in detecting DoS attacks. However, these methods typically require large amounts of labeled data for training, which may not be available in real-world scenarios, especially when timely response to novel or unseen attacks is required.
In addition to detection, several mitigation strategies have been proposed. For example, Gupta et al. [22] developed a machine learning (ML)-based Distributed DoS (DDoS) mitigation framework that leverages SDN’s programmability and centralized control to enable intelligent traffic management in UAV environments. Recent research has also investigated adaptive control strategies to enhance the resilience of UAV swarms against DoS attacks. Wu et al. [23] proposed a zero-sum differential game approach to effectively handle connectivity disruptions caused by DoS attacks and ensure boundedness and consensus control performance among UAVs. Tang et al. [24] incorporated robustness constraints to enhance disturbance resilience and introduced a dynamic event-triggered mechanism to respond to varying attack durations. Nevertheless, these approaches usually enhance the robustness of UAV swarms through tolerance and isolation strategies, which may be insufficient against sophisticated adversaries.
Compared with the above methods, our approach offers several significant advantages. First, our method does not rely on prior knowledge, labeled data, or attack signatures, making it more robust to novel and adaptive threats. Second, by orchestrating multiple lightweight MTD mechanisms, our framework can rapidly respond to attacks in real time, minimizing service disruption and resource consumption. Third, the federated multi-agent reinforcement learning design enables collaborative defense without sharing raw data, which is particularly suitable for resource-constrained environments. These features collectively distinguish our work from prior studies and address key challenges in securing low-altitude networks.
@@ II-B MTD-based Solutions for DoS/DDoS Attacks
The core idea of MTD is to reduce the predictability and exposure of critical system components by introducing dynamic changes to the system’s configuration or behavior. Therefore, typical MTD techniques are designed to invalidate the attacker’s reconnaissance efforts, force them to adapt continuously, and ultimately increase the likelihood of attack failure. In recent years, MTD has been widely studied to counter DoS/DDoS attacks in both traditional and emerging network environments [25, 26, 27].
For example, Ribeiro et al. [28] proposed a novel architecture leveraging Software Defined Networking (SDN) for flow classification and MTD techniques for DDoS mitigation. Similarly, Zhang et al. [29] developed a collaborative mutation-based MTD (CM-MTD) framework to disrupt attacks. The authors formulated the MTD deployment as a semi-Markov decision process (SMDP), leveraging a hierarchical deep reinforcement learning (DRL) algorithm for scheduling of MTD actions. However, these MTD mechanisms are equipped with high computation resources, which are not suitable for resource-constrained environments such as UAV networks.
To enable DoS/DDoS defense in resource-constrained environments, several recent studies have explored lightweight MTD mechanisms. For instance, Zhang et al. [30] designed configuration mutation mechanisms against DDoS attacks, where the communication ranges and capacities of roadside units (RSUs) are dynamically adjusted using a DRL algorithm. Similarly, Zhou et al. [31] introduced lightweight MTD mechanisms for the Internet of Things (IoT) environments, focusing on the dynamic control of device admission and service replica migration to defeat DDoS attacks. In the edge cloud context, the authors [18] designed several container-based MTD mechanisms and employed a deep Q-network (DQN) algorithm to optimize the trade-off between security and overhead when mitigating DDoS attacks. However, these mechanisms are not directly applicable to DoS defense in UAV networks. Furthermore, most existing studies that employ DRL for MTD optimization adopt centralized decision-making paradigms, which are ill-suited for the distributed and dynamic nature of UAV swarms.
In summary, while prior MTD-based solutions have demonstrated effectiveness in specific scenarios, our work is the first to propose a federated and multi-agent MTD framework specifically tailored for low-altitude networks. Unlike previous studies that rely on centralized decision-making or heavyweight mutations, our approach enables distributed, adaptive, and energy-efficient defense by leveraging lightweight mechanisms and federated learning. Furthermore, our method optimizes MTD deployment through intelligent policy learning, achieving superior robustness and scalability in dynamic, low-altitude network scenarios. These unique advantages highlight the novelty and practical value of our contribution.
@@ III-A System Model
The UAV swarm is required to maintain a circular formation of radius $\mathbf{r}$ centered at $\mathbf{c}=(c_{x},c_{y},h)$. The ideal position for UAV $i$ at time $t$ is given by
$\mathbf{p}_{i}^{*}(t)=\mathbf{c}+\mathbf{r}\cdot\begin{bmatrix}\cos\theta_{i}(t)\\ \sin\theta_{i}(t)\\ 0\end{bmatrix}$ (2)
@@
where $\theta_{i}(0)$ is the initial angular position of UAV $i$. The patrol (linear) speed of each UAV along the circle is given by $v_{\text{pat}}=r\omega$, where $r$ is the formation radius. To ensure collision avoidance, the UAVs are required to maintain a minimum distance $d_{\min}$ from each other at all times such that
@@ III-B2 Attacker Strategies
  • Greedy Attacker: The adversary dynamically observes the network state, enabling adaptively selecting the target that is expected to cause the maximum disruption to the swarm [36]. For example, the greedy attacker can select the current leader UAV $\mathbb{L}^{\mathcal{N}}(t)$ or the core link $\mathbb{L}^{\mathcal{L}}(t)$ of the swam as the target at time step $t$. For the sake of not weakening the greedy attacker’s ability, we assume that he/she may be aware of the existence of defense mechanisms (e.g., MTD), but not their specific deployment or timing. Therefore, the attacker can re-evaluate and update his/her target at time step $t^{\prime}$ to maximize attack effectiveness, but attack actions are subject to duration and reconnaissance constraints that are similar to those of the random attacker, i.e., $t^{\prime}=t+\tau_{\text{atk}}+\tau_{\text{recon}}$. This reconnaissance period represents the time required for the adversary to gather new intelligence, such as identifying the new leader after a switch or discovering the new communication channel after a frequency hop, before launching the next targeted attack.
@@ III-E2 Action Space
Specifically, $\tilde{A}_{t}^{\mathbb{L}}=\{\mathbf{l}_{1}(t),\mathbf{l}_{2}(t),...,\mathbf{l}_{N}(t)\}$ denotes the decisions of leader switching at time step $t$. Similarly, $\tilde{A}_{t}^{\mathbb{R}}=\{a_{t,1}^{\mathbb{R}},a_{t,2}^{\mathbb{R}},...,a_{t,N}^{\mathbb{R}}\}$ denotes the actions of route mutation for all UAVs, where $a_{t,n}^{\mathbb{R}}=1$ represents that the $i$-th UAV plays the role of relay node whereas $a_{t,n}^{\mathbb{R}}=-1$ means canceling its role. Although the action space is formally defined for all agents, the initiation of global MTD mechanism (i.e., frequency hopping) is a privileged operation executed only by the current leader agent $\mathbb{L}^{\mathcal{N}}(t)$. When the leader’s policy selects frequency hopping, the corresponding command is broadcast to all follower agents for synchronized execution. We define it as $\tilde{A}_{t}^{\mathbb{F}}=\{a_{t,1}^{\mathbb{F}},a_{t,2}^{\mathbb{F}},...,a_{t,N}^{\mathbb{F}}\}$, where $\tilde{A}_{t}^{\mathbb{F}}=[1,1,\ldots,1]$ triggers the change of the communication frequency for all UAVs and the GCS. This ensures that while all agents learn a comprehensive policy, only the leader’s decision on global matters is enacted, thus resolving the coordination challenge.
@@ III-E3 Transition Probability
Let $S_{t}\times A_{t}\times S_{t+1}\rightarrow[0,1]$ represent the state transition function. The environment transitions to a new state $s^{\prime}$ according to the transition probability $P(s^{\prime}|s,a)$, where $s$ is the current state at time $t$ and $a\in A_{t}$ is the action selected by all agents. This probability is influenced by the system state, the attacker’s strategy, and the defender’s chosen actions. In the case of deterministic transitions, the probability becomes binary and can be formulated as
@@
For instance, if the defender takes no action (i.e., all elements in $A_{t}$ are zero) on all UAVs and the attacker continues with the same malicious requests, the system remains in its current state, resulting in $P(s^{\prime}|s,a)=0$. Conversely, if the defender initiates a global frequency hopping when there is no ongoing attack, the resulting system state can be determined directly by this action.
@@ IV-B Local Policy Optimization
We select a policy gradient (PG) based approach for local policy optimization due to its specific advantages in addressing the UAV defense problem. First, PG methods naturally learn a stochastic policy, which aligns perfectly with the MTD principle by introducing unpredictability into the UAV swarm. This makes it significantly harder for an attacker to counter the defensive actions. Second, PG methods are well-suited for the complex, multi-dimensional action space of our problem and generally offer more stable convergence properties compared to value-based methods like DQN, which is crucial for learning a reliable policy in our dynamic environment.
In detail, deep neural network (DNN) models are utilized to build the learning agent in the loacl scenario. Each agent maintains a policy network $\pi_{\theta_{i}}(a|o)$ parameterized by $\theta_{i}$. At each time step $t$, the agent receives a local observation $o_{t}^{i}$ and samples an action $a_{t}^{i}$ from the policy as
@@ IV-C Federated Parameter Aggregation
After a number of local updates (e.g., $K$ episodes), all agents upload their local parameters $\theta_{i}$ to the central aggregator. Instead of naive averaging, we perform a reward-weighted aggregation of policy parameters. For each participating agent $i$, we compute a weight $w_{i}$ proportional to its recent average return over the most recent $M$ episodes, denoted as
@@ IV-D1 Algorithm Description
Specifically, at the beginning of the training process, we first initialize the global shared parameters $\theta^{\text{global}}$ in Line 1. For each agent $i$, the local policy parameters $\theta_{i}$ are set to $\theta^{\text{global}}$, and the local experience buffer $\mathcal{D}i$ is initialized as empty in Lines 2 and 3, respectively. Subsequently, Line 4 starts the main training loop, which iterates for a total of $K_{\max}$ episodes.
Within each episode, as shown in Lines 5–12, each agent $i$ resets its environment and observes the initial state. At each time step $t$, the agent selects an action $a_{t}^{i}$ according to its current policy $\pi_{\theta_{i}}$, receives the immediate reward $r_{t}^{i}$ and observes the next state $o_{t+1}^{i}$. The transition tuple $(o_{t}^{i},a_{t}^{i},r_{t}^{i},o_{t+1}^{i})$ is then stored in the local experience buffer $\mathcal{D}_{i}$. After completing the episode, it computes the return $\Omega_{t}^{i}$ and normalized return $\hat{R}_{t}^{i}$, and subsequently updates $\theta_{i}$ by minimizing the loss function $L(\theta_{i})$ using the policy gradient method.
Every $K$ episodes, as indicated in Lines 13–21, all agents participate in the aggregation process. Each agent computes its average return $\bar{R}i^{(M)}$ over the recent window of $M$ episodes and uploads both its shared parameters and average return to the central aggregator. The aggregator then calculates the reward-weighted aggregation coefficients $w_{i}$ and aggregates the shared parameters to obtain the updated global parameters $\theta^{\text{global}}$. After that, each agent synchronizes its shared layers and performs $T{\text{local}}$ steps of local fine-tuning, further updating its policy parameters to adapt to local environments.
1 Initialize global shared parameters $\theta^{\text{global}}$ ;
@@ 18         Upload $\theta_{i}$ and $\bar{R}_{i}^{(M)}$ to the aggregator;
20     Compute weights $w_{i}$ (Eq. (29)) and aggregates parameters using Eq. (30);
@@ IV-D2 Complexity Analysis
In each episode, the agent interacts with the environment for $T$ time steps, leading to a per-episode complexity of $O(T\sum_{n=1}^{L_{\text{P}}-1}N_{n}N_{n+1})$. After every $K$ episodes, the federated aggregation step involves computing the weighted average of the shared parameters across $N$ agents, which has a complexity of $O(NP)$, where $P$ is the number of shared parameters. The local fine-tuning phase after aggregation consists of $T_{\text{local}}$ gradient updates per agent, each with complexity $O(\sum_{n=1}^{L_{\text{P}}-1}N_{n}N_{n+1})$. Therefore, the overall computational complexity of the proposed method is $O(NKT\sum_{n=1}^{L_{\text{P}}-1}N_{n}N_{n+1}+NP+NT_{\text{local}}\sum_{n=1}^{L_{\text{P}}-1}N_{n}N_{n+1})$.

### v4

@@ I Introduction
Nevertheless, the widespread adoption of UAV swarms also brings new security challenges [7, 8] to low-altitude networks. Due to their reliance on open wireless links, limited energy and processing capabilities, UAV networks are particularly vulnerable to Denial-of-Service (DoS) attacks [9, 10]. For example, UAVs often operate without robust authentication or traffic filtering mechanisms. Attackers can easily launch attacks by overwhelming low-altitude communication channels or computational resources, leading to service disruptions or UAV disconnection from the swarm [11]. In mission-critical scenarios, even a brief loss of connectivity or control can have catastrophic consequences, including the failure of time-sensitive missions or the crash of drones.
@@ II-A DoS Attacks Detection and Mitigation in UAV Networks
Research on securing UAV networks against DoS attacks has largely focused on two areas, including detection and mitigation. Detection methods often employ machine learning models like convolutional neural network (CNN) [20], long short-term memory (LSTM) [21] or conditional generative adversarial network (CGAN) [22] to identify malicious traffic. While achieving high accuracy, these approaches rely heavily on large, labeled datasets, making them less effective against novel or zero-day attacks and difficult to deploy in real-world scenarios where such data is scarce.
In addition to detection, mitigation strategies often involve adaptive control or game-theoretic approaches to enhance swarm resilience [23]. For example, Tan et al. [24] proposed a time-delay evolutionary game method for honeypoint defense, ensuring optimal strategy stability and selects the best defense at the 35th time unit on IEEE 22-Bus. These methods typically focus on tolerating or isolating disruptions rather than proactively preventing them. A key limitation of such reactive strategies is that they often act only after an attack has already caused damage, which may be insufficient against sophisticated adversaries who can adapt their tactics to circumvent static or predictable defenses.
Compared with the above methods, our approach offers two significant advantages. First, our method does not rely on prior knowledge, labeled data, or attack signatures, making it more robust to novel and adaptive threats. Second, by orchestrating multiple lightweight MTD mechanisms, our framework can rapidly respond to attacks in real time, minimizing service disruption and resource consumption. These features collectively distinguish our work from prior studies and address key challenges in securing low-altitude networks.
@@ II-B MTD-based Solutions for DoS/DDoS Attacks
MTD has emerged as a proactive paradigm to counter DoS or DDoS attacks by dynamically altering the system’s attack surface [25, 26]. In traditional scenarios, MTD often leverages software-defined networking (SDN) or virtualization to shuffle IP addresses or migrate services [27, 28]. However, these solutions present a significant limitation: their high computational and communication overhead makes them unsuitable for resource-constrained environments like low-altitude networks.
To address this, some research has explored lightweight MTD methods for Internet of Things (IoT) and edge environments, using deep reinforcement learning (DRL) to optimize configuration changes [29, 19]. For instance, Zhang et al. [30] designed configuration mutation mechanisms against DDoS attacks, where the communication ranges and capacities of roadside units (RSUs) are dynamically adjusted. While more efficient, these methods face another key limitation. They are not directly applicable to the highly distributed and mobile nature of UAV swarms, where centralized decision-making creates bottlenecks and single points of failure.
In contrast to these prior works, our framework is the first to integrate lightweight, coordinated MTD mechanisms with a federated multi-agent reinforcement learning approach. This design specifically addresses the current limitations of existing MTD solutions by enabling proactive, distributed, and intelligent defense tailored for the unique challenges of low-altitude networks.
@@ III-A System Model
The UAV swarm is required to maintain a circular formation of radius $r$ centered at $\mathbf{c}=(c_{x},c_{y},h)$. The ideal position for UAV $i$ at time $t$ is given by
$\color[rgb]{0,0,0}\mathbf{p}_{i}^{*}(t)=\mathbf{c}+r\begin{bmatrix}\cos\theta_{i}(t)\\ \sin\theta_{i}(t)\\ 0\end{bmatrix}$}} (2)
@@
where $\theta_{i}(0)$ is the initial angular position of UAV $i$. The patrol (linear) speed of each UAV along the circle is given by $v_{\text{pat}}=r\omega$. To ensure collision avoidance, the UAVs are required to maintain a minimum distance $d_{\min}$ from each other at all times such that
@@ III-B2 Attacker Strategies
  • Greedy Attacker: The adversary dynamically observes the network state, enabling adaptively selecting the target that is expected to cause the maximum disruption to the swarm [35]. For example, the greedy attacker can select the current leader UAV $\mathbb{L}^{\mathcal{N}}(t)$ or the core link $\mathbb{L}^{\mathcal{L}}(t)$ of the swarm as the target at time step $t$. For the sake of not weakening the greedy attacker’s ability, we assume that he/she may be aware of the existence of defense mechanisms (e.g., MTD), but not their specific deployment or timing. Therefore, the attacker can re-evaluate and update his/her target at time step $t^{\prime}$ to maximize attack effectiveness, but attack actions are subject to duration and reconnaissance constraints that are similar to those of the random attacker, i.e., $t^{\prime}=t+\tau_{\text{atk}}+\tau_{\text{recon}}$. This reconnaissance period represents the time required for the adversary to gather new intelligence, such as identifying the new leader after a switch or discovering the new communication channel after a frequency hop, before launching the next targeted attack.
@@ III-E2 Action Space
Specifically, $\tilde{A}_{t}^{\mathbb{L}}=\{\mathbf{l}_{1}(t),\mathbf{l}_{2}(t),...,\mathbf{l}_{N}(t)\}$ denotes the action of leader switching at time step $t$. If a switch is performed, exactly one element $\mathbf{l}_{i}(t)$ will be 1, designating UAV $i$ as the new leader. If no switch occurs, all elements of $\tilde{A}_{t}^{\mathbb{L}}$ will be 0. Similarly, $\tilde{A}_{t}^{\mathbb{R}}=\{a_{t,1}^{\mathbb{R}},a_{t,2}^{\mathbb{R}},...,a_{t,N}^{\mathbb{R}}\}$ denotes the actions of route mutation for all UAVs, where $a_{t,n}^{\mathbb{R}}=1$ represents that the $i$-th UAV plays the role of relay node whereas $a_{t,n}^{\mathbb{R}}=-1$ means canceling its role. Although the action space is formally defined for all agents, the initiation of global MTD mechanism (i.e., frequency hopping) is a privileged operation executed only by the current leader agent $\mathbb{L}^{\mathcal{N}}(t)$. When the leader’s policy selects frequency hopping, the corresponding command is broadcast to all follower agents for synchronized execution. We define it as $\tilde{A}_{t}^{\mathbb{F}}=\{a_{t,1}^{\mathbb{F}},a_{t,2}^{\mathbb{F}},...,a_{t,N}^{\mathbb{F}}\}$, where $\tilde{A}_{t}^{\mathbb{F}}=[1,1,\ldots,1]$ triggers the change of the communication frequency for all UAVs and the GCS. This ensures that while all agents learn a comprehensive policy, only the leader’s decision on global matters is enacted, thus resolving the coordination challenge.
@@ III-E3 Transition Probability
The state transition function $P(S_{t+1}|S_{t},A_{t})$ defines the probability of transitioning to state $S_{t+1}$ given the current state $S_{t}$ and the joint action $A_{t}$. In our simulated environment, the system dynamics are deterministic. This means that for any given state and action pair, the next state is uniquely determined. Formally, the transition probability can be expressed as
@@
For instance, if the defender takes no action (i.e., all elements in $A_{t}$ are zero) on all UAVs and the attacker continues with the same malicious requests, the system remains in its current state, i.e., $P(s^{\prime}|s,a)=1$ when $s^{\prime}=s$ (and 0 otherwise). Conversely, if the defender initiates a global frequency hopping when there is no ongoing attack, the frequency components of the next state are directly determined by this action, resulting in a predictable state transition with $P(s^{\prime}|s,a)=1$ for a specific next state $s^{\prime}$.
@@ IV-B Local Policy Optimization
We select a policy gradient (PG) based approach for local policy optimization due to its specific advantages in addressing the UAV defense problem. First, PG methods naturally learn a stochastic policy, which aligns perfectly with the MTD principle by introducing unpredictability into the UAV swarm. This makes it significantly harder for an attacker to counter the defensive actions. Second, PG methods are well-suited for the complex, multi-dimensional action space of our problem and generally offer more stable convergence properties compared to value-based methods like deep Q-network (DQN), which is crucial for learning a reliable policy in our dynamic environment.
In detail, deep neural network (DNN) models are utilized to build the learning agent in the local scenario. Each agent maintains a policy network $\pi_{\theta_{i}}(a|o)$ parameterized by $\theta_{i}$. At each time step $t$, the agent receives a local observation $o_{t}^{i}$ and samples an action $a_{t}^{i}$ from the policy as
@@ IV-C Federated Parameter Aggregation
After a number of local updates (e.g., $K_{\text{agg}}$ episodes), all agents upload their local parameters $\theta_{i}$ to the central aggregator. Instead of naive averaging, we perform a reward-weighted aggregation of policy parameters. For each participating agent $i$, we compute a weight $w_{i}$ proportional to its recent average return over the most recent $M$ episodes, denoted as
@@ IV-D1 Algorithm Description
Algorithm 1 details the training process for our PG-FMADRL method. The procedure iterates for $K_{\max}$ episodes, alternating between local policy updates and periodic federated aggregation. In the local phase, each agent $i$ interacts with the environment, collecting experience tuples $(o_{t}^{i},a_{t}^{i},r_{t}^{i},o_{t+1}^{i})$ into its buffer $\mathcal{D}_{i}$. At the end of each episode, it updates its local policy parameters $\theta_{i}$ by minimizing the loss function $L(\theta_{i})$ (Eq. (27)) via policy gradient.
The federated aggregation phase commences every $K_{\text{agg}}$ episodes, during which each agent contributes its local parameters $\theta_{i}$ and recent average reward $\bar{R}_{i}^{(M)}$ to the central aggregator. The aggregator then computes an improved global model $\theta^{\text{global}}$ using reward-weighted averaging (Eq. (30)). Finally, agents synchronize their shared layers with $\theta^{\text{global}}$ and perform $T_{\text{local}}$ local fine-tuning steps to adapt the generalized policy to their specific context.
1Initialize global shared parameters $\theta^{\text{global}}$ ;
@@ 18         Upload $\theta_{i}$ and $\bar{R}_{i}^{(M)}$ to the aggregator;
20     Compute weights $w_{i}$ (Eq. (29)) and aggregate parameters using Eq. (30);
@@ IV-D2 Complexity Analysis
In each episode, the agent interacts with the environment for $T$ time steps, leading to a per-episode complexity of $O(T\sum_{n=1}^{L_{\text{P}}-1}N_{n}N_{n+1})$. After every $K_{\text{agg}}$ episodes, the federated aggregation step computes the weighted average of $P$ shared parameters across $N$ agents, which has a complexity of $O(NP)$. Then, the local fine-tuning phase consists of $T_{\text{local}}$ gradient updates per agent, each with complexity $O(\sum_{n=1}^{L_{\text{P}}-1}N_{n}N_{n+1})$. Thus, the overall computational complexity is $O(NK_{\text{agg}}T\sum_{n=1}^{L_{\text{P}}-1}N_{n}N_{n+1}+NP+NT_{\text{local}}\sum_{n=1}^{L_{\text{P}}-1}N_{n}N_{n+1})$.
@@ V Performance Evaluation
V-A Simulation Setup
The reward coefficients $(\alpha,\beta,\zeta,\eta,\xi)$ and entropy coefficient $\mu$ were determined through an iterative tuning process. Our methodology prioritized critical objectives, setting a high weight for the attack penalty $\eta=2$ to focus on threat mitigation, followed by the velocity penalty $\xi=0.5$ to limit energy use. The remaining coefficients were fine-tuned around $[0,1]$. Each configuration was evaluated based on learning stability, final converged reward, and attack mitigation rate. The final values for all simulation parameters are summarized in Table I with reference to [37, 38, 39].